Home‎ > ‎Reports‎ > ‎Crew097‎ > ‎

day08


Crew 97 Log Book for January 8, 2011

Photos of the day





Commander's Check-In Report

Judah Epstein Reporting

Crew Physical Status:

Good.


Time Departed/Returned from EVA:

EVA #11: 10:16 / 11:00.

EVA #12: 12:18 / 13:22.

EVA #13: 14:22 / 16:21. Sandstorm Brief Narrative of Field Mission Results:

EVAs today consisted of measuring water gully slush flows, locating rock varnish, and operating the Sandstorm Rover.


EVA Data/Interpretations:

EVA #11: Two water gully slush flows located and elevation levels taken.

EVA #12: Two locations with rock varnish were located.

EVA #13: Tested travel distance capability of the rover.


Engineering/Hab Maintenance:

n/a

Report Transmission Schedule:

CMDR Check-in, ENGR, CMDR Report, Photos, 3 EVA Reports, 3 Science Reports


Plans for Tomorrow:

- Rest day
- Sleep in
- Breakfast
- Psychophysiological Experiment
- EVA Geocache
- Lunch
- EVA exploration
- Dinner
- Reports by 20:00
- Food Study Questionnaire
- Showers


Inventory:

n/a

Support Requested:

n/a

Miscellaneous:

DG took control panel from water heater to replace components.

Commander's Report

Judah Epstein Reporting

Marsonauts are successfully progressing with today's multiple experiments and EVAs.

The morning EVA consisted of measuring water gully slush flows for the Crew Geologist's experiment. Experiment preparation also consisted of preparation of the rover as well as psycho-physiological tests on crew-members in which each crew-member was exposed to digital images and sounds and affective response was measured through heart rate variability and skin conductance level. The Commander's Water Filtration systems were analyzed with additional measurements and bacteria results.

The afternoon consisted of two EVAs. First was a reconnaissance EVA whereas rock varnish areas were successfully identified. The third EVA of the day consisted of testing travel distance capability of the rover.

Although tomorrow is a scheduled rest day, the crew is eager to conduct additional EVAs and continue progress on respective science experiments.

Engineering Report

Nathan Wong Reporting

Generator/Electricity:

Kitty(Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level:

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off): Power went off once today.


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): .75

Propane Tank (Full/0.75/0.5/0.25/Empty): .55

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 3 Gas cans full

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): 3

Honey Oil Quantity (# of Quarts in storage): 2

ATV Oil Quantity (# of Quarts in storage): 2

Car Oil Quantity (# of Quarts in storage): none


Water Status:

Outside Potable Water Tank Level (inches from bottom): half%

Trailer Potable Water Tank Level (inches from bottom): full

Hab Potable Water Tank Level (inches from bottom): Full%

Potable Water Meter Reading: 40672.5

In to GreenHab Meter Reading:

Notes/Comments:


GreenHab:

Condition of plants in Tank 1: None

Condition of plants in Tank 2: 3 plants

Condition of plants in Tank 3: 7 plants

Greenhab Notes/Comments: Plants in tank two are not floating upright. I will take a picture and send it with tomorrow's report.


Observatory:

Telescope Used? (Yes/No): No

Observatory Notes/Comments: Observatory not used


Transportation:

Hab Car (NOT RENTAL)
Used (yes/no): No
Oil Checked (yes/no): No
Tire Status: good
Odometer at end of day: 123085
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: Yes
Oil Checked: No
Fuel Consumed: No
Tire Status: Inflated
Comments:

ATV 2 (Spirit):
Used: Yes
Oil Checked: No
Fuel Consumed: No
Tire Status: Inflated
Comments:

ATV 3 (Viking I):
Used: No
Oil Checked: No
Fuel Consumed: No
Tire Status: Inflated
Comments:

ATV 4 (Viking II):
Used: Yes
Oil Checked: No
Fuel Consumed: Yes
Tire Status: Inflated
Comments:


Heating and Ventilation:
Thermostat setting upstairs (Farenheit): off
Actual temperature upstairs (Farenheit): 74

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 57


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: Everything is working well


General Engineering Notes/Comments:

Toilet is slow to flush. I will use the toilet snake to see if that helps.

Science Report 1

Judah Epstein Reporting

The purpose of this project is to explore the effectiveness of a slow sand sedimentary water filtration system in a planetary environment. To simulate a planetary environment the experiment is conducted at the Mars Desert Research Station, a self-sustaining planetary research experiment and simulation expedition.

Background information regarding the project abstract, purpose, goals, and specific experiment details and information are included in the Science Report from January 6.

___________________________________

The following data was collected yesterday (January 7):

MISSION SUPPORT NOTE: Spreadsheet of the Jan 7th data is here.


Science Report 2

Tonya Thompson Reporting

I performed a psychophysiological test of affective response to digital images and sounds on each crewmember 2 times today. The images were projected onto a simulated 3D surface and the sounds were played through headphones.

Each crewmember sat in the test area in my stateroom for 11 minutes (a 3 minute pre-exposure baseline, a 5 minute affective images and sounds exposure, and a 3-minute post exposure measurement). During this time, I measured the crewmember’s heart rate variability and skin conductance level in response to the digital images and sounds.

I also had each member complete a Cohen’s Perceived Stress Scale questionnaire after dinner.

Science Report 3

Lucinda Land Reporting

Day two of testing Sandstorm's range was a success. Yesterday we lost touch with Sandstorm at 193.12 meters outside of the Hab with the router inside the lab area. We were testing the range of the router signal while inside the Hab with the rover. Sandstorm went an additional 16.09 meters today to make her total distance with the router inside the Hab 209.21 meters.

We put fresh batteries in Sandstorm and then took her out to the same place yesterday where she stopped. We wanted to confirm if it was the range of the signal or dead batteries for the reason that she no longer responded. Once the XO and the rover finished pressurization in the airlock he hand carried her to the GPS location he logged yesterday as her stopping point. We still had connection to Sandstorm at that point and drove her a further 16 meters. The obvious issues with long range driving of the rover was video confirmation as the image on the laptop screen would freeze. All the software that drives the rover is created by our NorCal Mars Society team and one idea that came forward in today's test was to have a refresh button built into the rover software to refresh her camera to hopefully unfreeze transmission when operating. Reducing frames per second (fps) to her lowest setting didn't make any significant changes with video feed.

Another issue with wi-fi signal to connect with Sandstorm was line of site and the presence of hilly terrain. Sandstorm had a bit of trouble traversing deep snow and steeper mounds. We tried to keep her on as flat a surface as possible. She could however make it through small dips and bumps in the road with an increase of the throttle. Larger rocks would often impede her movement but the Marsonauts would aid her in maneuvering around it.

One example of better maneuvering technique with the aid of a Marsonaut was when the Crew Engineer was visible in her forward camera view. This way Sandstorm was able to follow in a relatively clear path. This was most helpful when Sandstorm experienced very rocky terrain and Crew Engineer Wong walked in front of Sandstorm to lead her through without incident This will be a future technique to use when taking her out on another remote exploration.

There was a change of batteries made during her nearly two hour EVA with the XO later joined by the Crew Engineer. Since Sandstorm's batteries only last an hour, we had to change them out for her to continue her tests. This was done with the use of the Hab's ATV pressurized garage. During the second half of her EVA, Sandstorm's router was moved outside the Hab to test her range. Sandstorm logged 128.74 meters with the router outside the Hab. Her batteries died prematurely and she was rendered temporarily inoperable. We made sure to determine that her batteries were dead by removing any other possible variables such as hilly terrain and distance. The XO brought her back into the router's line of sight and on flatter terrain where operation of her was still not possible.

We will start Sandstorm up again tomorrow and give her fresh batteries. So far so good on getting her started up and in motion. We have been quite successful with getting her going. We just wish she could go longer and that the signal strength was strong enough to deliver adequate video feed to confirm her placement and direction of movement. Using her in conjunction with Marsonauts proved helpful when we couldn't rely on her camera for confirmation of placement, direction, or terrain.

Tomorrow we will take her outside for a second test with the router outside and using fresh batteries. We will measure a second distance and see if we can beat the 128.74 meters she logged in this configuration. My hypothesis was that she would go for a longer distance with the router outside the Hab than inside. However, our two days of testing with the router inside proved that she can go a distance of at least 209.21 meters. Now we need to prove whether she gets more distance or not with the router outside the Hab.

In her near future we will continue tests with the use of her Gigapan camera and GPS logging systems.

EVA 11 Report

Jim Crowell Reporting

Time: 10:16 / 11:00

Crew Members: Amanda Damptz, Jim Crowell

Site Location: North of the Hab, along the road

Transit Mode: ATVs

Objectives: To locate and measure slush flow locations

Results: Two slush flows located and elevation levels taken

Lessons Learned: Defogging the facemask, sealing the helmet, and push-to-talk radio communication are key things to check before a ATV EVAs.

EVA 12 Report

Jim Crowell Reporting

Time: 12:18 / 13:22

Crew Members: Nathan Wong, Judah Epstein, Jim Crowell

Site Location: West of the Hab, along the ATV path, accessed North of the Hab

Transit Mode: ATVs

Objectives: Locate areas rich with rock varnish

Results: Two locations found. Samples to be taken later in the week.

Lessons Learned: All locations did not need up-close analysis; varnish could be seen from a distance.

EVA 13 Report

Jim Crowell Reporting

Time: 14:22 / 16:21

Crew Members:

Site Location: Near the Hab, toward the East

Transit Mode: Walking

Objectives: To test if the rover known as Sandstorm could travel more than 193.12 metres away from the Hab

Results: Sandstorm was walked out to the location 193.12 metres away from the Hab. Several attempts were made by controller (- inside the Hab), and the rover was moved a few metres; when it would no longer respond, it was carried halfway back and controlling was attempted again with no success. It was determined the batteries were low, so the rover was placed in the ATV garage. Once the garage repressurized, entered and replaced the batteries. She went back into the Hab, and the garage was depressurized so Jim could receive the rover again. At this time, J began to experience difficulties with his suit's respiratory system, so N. suited up to assist. The garage was then repressurized so L. could set up control station in the ATV garage. 2 crew then followed the rover as L. controlled it. When the batteries died again, the rover was carried back to the Hab.

Lessons Learned: Rover can function farther than before, if batteries are charged and there is a clear view of the controller. Respiratory systems should be double checked prior to EVA.

Cook's Report

Nathan Wong Reporting

1a) Type of day:
-A cooking day? Yes
-Celebration foods? No

2) List the foods served at today's main meal
Pesto Pasta
Cheesy Beef Potatoes

3) List any main meal foods not finished by the crew
None

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

Cheesy Beef Potatoes
1. Rehydrate freeze dried beef burger crumbles
2. Make Au Gratin Potato Package
3. Combine
Pesto pasta
1. Cook pasta
2. Prepare instant pesto sauce and creamy pesto sauce together according to package directions
3. Combine

5) Comments/questions
None